# MSCCL++: Rethinking GPU Communication Abstractions for Cutting-Edge AI Applications

### v3

@@ 3.1. Hierarchical User Interfaces
MSCCL++ Primitive API. This replaces NCCL’s primitives with a set of core communication functions that are called from the GPU kernels. This is paired with the bootstrapping API called from the host (CPU) side, which is used for connection setup between GPUs. The primitive interface is the base layer that is used to implement the other two higher-level interfaces. Programmers can also use the primitive interface directly to implement application-specific or hardware-specific optimizations in their own GPU kernels, including those kernels that also perform computation. As the primitive interface is our key contribution, we focus on its details in Section 3 and 4.
@@
MSCCL++ DSL API. We reimplement the MSCCLang (Cowan et al., 2023) DSL over MSCCL++, which allows us to write custom collective communication algorithms in a high-level language. This interface is aimed at users who want to specify communication algorithms or schedules and generate kernels that are optimized for their own workloads (application and hardware). The generated kernel uses the Primitive API by running the DSL Executor, so it may have small runtime overheads.
@@
MSCCL++ Collective API. We reimplement the NCCL API, including its bootstrapping API, as-is over MSCCL++. It is aimed at users with the least expertise; they can simply replace NCCL/RCCL with the MSCCL++ Collective library without changing their application code. This library has the same limitations as that of NCCL (Section 2.3); i.e., it may not provide the best algorithm for certain workloads. For AI inference workloads, we implement the best algorithms in our collective kernels using the MSCCL++ Primitive API.
3.2. Primitive Communication Abstractions
3.2.1. Communication Channels
@@
The supported mode(s) of data transfer differs between interconnects. Therefore, MSCCL++ defines an abstract channel for each data transfer mode supported by the hardware, and exposes communication primitives specific to that channel that can be invoked directly inside a GPU kernel. MSCCL++ defines three communication channels (for readers who are familiar with NCCL, MSCCL++ channel does not correspond to that in NCCL): PortChannel, MemoryChannel, and SwitchChannel, which correspond to port-mapped I/O, memory-mapped I/O, and switch-mapped I/O respectively. A PortChannel uses interconnect ports to communicate between GPUs; i.e., a GPU can initiate the data transfer through a port to another GPU. Ports are controlled by dedicated hardware for I/O, such as DMA engines on GPUs or RDMA NICs. A MemoryChannel uses peer-to-peer memory access to communicate between GPUs; i.e., a GPU can directly access another GPU’s memory. A SwitchChannel uses interconnection-switch-enabled multimem memory access to communicate between GPUs; i.e., a GPU can access the memory of multiple GPUs simultaneously through a switch.
The same interconnect can support multiple modes of data transfer, so multiple channels can be supported on the same interconnect; e.g., all three channels are supported by NVLink, only PortChannel and MemoryChannel are supported by xGMI and PCIe, while only PortChannel is supported by InfiniBand. The implementation of each channel is specific to the interconnect. The SwitchChannel uses NVSwitch multimem instructions that support computation-on-switch. The MemoryChannel uses thread-copy data transfer mode over NVLink, xGMI, and PCIe. On the other hand, the PortChannel uses the DMA-copy data transfer mode over NVLink, xGMI, PCIe, and Infiniband. Current interconnects require a CPU thread to initiate the data transfer; e.g., transferring data over InfiniBand requires the CPU to call ibv_post_send which initiates the RDMA transfer from the sending GPU memory to the receiving GPU memory. Such details are handled transparently by the implementation.
@@
Generality. Port-mapped I/O, memory-mapped I/O, and switch-mapped I/O are complementary I/O methods in general computer architectures. PortChannel, MemoryChannel, and SwitchChannel are our abstractions to provide these I/O methods from inside GPU kernels. We believe that they are flexible enough to accommodate future hardware advances. For example, if future hardware interconnects supports initiating DMA-copy from within a GPU kernel, the same PortChannel API can target the new hardware.
3.2.2. Communication Primitives
All MSCCL++ primitives are defined as a method of a channel. The primitives are all conceptual and do not imply any specific implementation. In this section, we describe the primitives API for the PortChannel with an example. Primitives for the other channels, and the implementation for all the channels are explained in Section 4.
Borrowing the term from MPI (Snir et al., 1998), writing data to the peer’s side through a channel is called put. To provide our communication abstraction as close as possible to hardware capabilities, we design put to be zero-copy (i.e., no intermediate buffers), one-sided (i.e., initiated by a peer without participation of the other peer), and asynchronous. MSCCL++ provides other interfaces for synchronization purposes. Specifically, signal and wait provide a mechanism for synchronization (including memory consistency) across GPUs. flush is for local synchronization that ensures all previous put operations are already on the fly.
Figure 4 shows the semantics of four MSCCL++ primitives: put, signal, wait, and flush. The put primitive transfers data from src0 buffer of GPU-0 to dst1 buffer of GPU-1. GPU-0 calls a following signal primitive that signals GPU-1, which is asynchronous yet strictly ordered with the previous put. GPU-1 has to call the wait primitive before it can read dst1. GPU-0 calls the flush primitive to ensure that the previous put is on the fly and src0 can be safely reused.
@@ 3.2.2. Communication Primitives
MSCCL++ also supports a few fused primitives that can reduce the overhead of API calls. For example, since a put is usually followed by a signal, we provide the putWithSignal function that conducts both at once.
1//async
2put(src0, dst1, size)
3//unsafe to reuse src0
4signal() //async
5flush() //sync
6//safe to reuse src0
@@
Example. Figure 5 shows MSCCL++ implementation of the all-pairs ReduceScatter algorithm, which is not used in NCCL. In this algorithm, all GPUs send their data at $\frac{i}{N}$ offset to GPU $i$, the $i^{th}$ GPU does elementwise reduction of the received data, and stores the reduced data. To store the received data, the implementation allocates a scratch buffer per GPU. The kernel’s main loop performs reduction in batches of the scratch size (line 518). On each GPU, the loop first put $\frac{1}{N}$ of data to all other GPUs’ scratch buffer and then signals the completion of transfer (line 910). Then a GPU waits for the data to arrive from other GPUs and reduce it with its part of input data (line 1315). Finally, a barrier among all GPUs is required to ensure scratch is not overwritten for next main loop offset (line 18). Section 5 shows that this algorithm works better than the ring algorithm for smaller data sizes while the ring algorithm works better for large data sizes.
3.2.3. Advantages of Communication Primitives
1global allPairsRS(count, gpus, channels[gpus])
2 sz = channel.scratchSz/gpus.num
3 count = count/gpus.num
4
5 for (off = 0; off < count; off += sz)
6 //Send 1/Nth data to each GPU
7 for (g = 0; g < gpus.num; g++)
8 idx = off + g * count;
9 channels[g].put(idx, sz*g, sz)
10 channels[g].signal()
11
12 //Reduce every pair of GPU
13 for (g = 0; g < gpus.num-1; g++)
14 channels[g].wait()
15 reduce(in + off, channels[g].scratch)
16
17 //barrier on all gpus
18 multiDeviceBarrier();
Figure 5. All-pairs ReduceScatter kernel in MSCCL++. Channels are initialized with source as input and destination as scratch buffer.
The MSCCL++ Primitive API itself does not introduce any novel optimization techniques – instead, it offers new straightforward abstractions that enable quick optimization of GPU communication. The key advantages come from exposing the primitive functionalities as a user interface. First, it allows application developers to easily and quickly optimize GPU communication algorithms for their own infrastructure and workloads. This is valuable in practice because the library developers may not be aware of the specific workloads and hardware configurations of their users. Second, the library developers can quickly deliver new hardware support by implementing the same primitive API for the new hardware. Third, since the primitive API is close to the hardware, it is more flexible to extend or adopt it for new hardware features.
MSCCL++ carefully optimizes GPU communication in the DSL Executor and collective kernels leveraging some of the below features in the Primitive API. Users of the Primitive API can also implement these optimizations directly.
Asynchronous Communication. MSCCL++ contains separate primitives for data transfer and synchronization, allowing asynchronous communication and batching synchronization of multiple data transfers. This abstraction has multiple benefits: (i) allows more algorithms to be specified than the synchronous communication in NCCL, (ii) enables faster collective communication performance by enabling algorithms to combine several synchronizations into one, (iii) frees up resources of peer GPUs for other meaningful work of the application, and (iv) enables better inter-kernel optimizations.
Specialized GPU Kernels. By using channels for specific GPU interconnects, MSCCL++ allows users to specialize kernels for particular interconnects. Furthermore, MSCCL++ avoid extra copies in kernels, leading to a kernel with less code paths and zero register spills. For example, for eight A100 GPUs, the ring AllReduce of NCCL and MSCCL use 94 and 96 registers per thread, respectively, while an MSCCL++ version uses only 32 per thread. Due to these reasons, MSCCL++ kernels execute less instructions, avoid memory accesses due to register spills, and have better instruction cache hits.
Interconnect Optimizations. NCCL only allows a single mode for data transfer over NVLink at a time. In contrast, by using different channel types in MSCCL++ (PortChannel or MemoryChannel), user can utilize both thread-copy and DMA-copy data transfer mode over NVLink.
1class Channel
2 protected:
3 //Src/Dst Buffers set during initialization
4 void* src, *dst;
5 //Semaphore alloced on this GPU
6 uint *semaphore; uint expectedVal;
7 public:
8 //Primitives
9 void signal() ; void wait()
10
11class PortChannel : public Channel
12 public:
13 void put(ulong dstOff, ulong srcOff, ulong sz,
14 uint tid, uint tids)
15
16template<Protocol protocol>
17class MemoryChannel : public Channel
18 public:
19 void put(ulong dstOff, ulong srcOff, ulong sz,
20 uint tid, uint tids, uint flag=0)
21 template<typename T>
22 T read(ulong off, uint tid, uint flag=0)
23
24 template<typename T>
25 void write(ulong off, uint tid, T elem,
26 uint flag=0)
27
28class SwitchChannel : public Channel
29 public:
30 void reduce(ulong dstOff, ulong srcOff,
31 ulong sz, uint tid, uint tids)
32 void broadcast(ulong dstOff, ulong srcOff,
33 ulong sz, uint tid, uint tids)
Figure 6. PortChannel, MemoryChannel, and SwitchChannel primitives in MSCCL++. Implementation of MemoryChannel primitives differ for the type of protocol: LL or HB.
4. MSCCL++ Design and Implementation
4.1. Initialization of MSCCL++
In general, a MSCCL++ program runs as a multi-process on all nodes and GPUs within each node, so that each process runs one GPU. MSCCL++ provides host (CPU) side interfaces for initializing multiple processes, and the interfaces can be specialized by users for their application. The initialization involves setting up connections between GPUs, finding the topology of GPU connections, and constructing channels and buffers.
The first step in initialization is to exchange metadata between host processes, which we call a Bootstrap. It consists of four virtual methods: (i) send to send metadata to another process, (ii) recv to receive metadata from a process, (iii) allGather to collect metadata of all processes, and (iv) barrier to synchronize all processes. MSCCL++ implements a default bootstrap methods using POSIX sockets, while it is straightforward to override with other popular implementations such as MPI or torch.distributed.
The second step is to construct a Communicator that registers communication buffers (either user buffers or dedicated ones) and creates corresponding MSCCL++ channels between GPUs based on the underlying physical links.
4.2. MSCCL++ Primitive API and Channels
A MSCCL++ channel is a connection between two or more GPUs. PortChannel and MemoryChannel are peer-to-peer (i.e., between two GPUs), whereas SwitchChannel is for a group of (i.e., two or more) GPUs. Figure 6 shows three types of channels and its interface in MSCCL++. A channel is initialized with the source and destination buffers, and an integer allocated on the GPU memory that servers as a semaphore.
4.2.1. PortChannel
A PortChannel implements primitives when data transfer is done over ports connected to GPU memory, such as cudaMemcpy for intra-node DMA or ibv_post_send for RDMA. Since data transfer over a port currently requires the CPU to initiate the transfer, each channel creates its own CPU thread. The CPU thread reads data transfer or synchronization requests from a first-in-first-out request queue. The storage, head, and tail of a request queue are allocated using cudaMallocManaged as both CPU and GPU can access it.
We now discuss the workflow of PortChannels using Figure 7. 0 When the GPU calls a primitive, such as put, the first participating thread of the GPU pushes the request to the queue by writing at the head. Before writing to the queue, the GPU checks if the queue is filled, i.e., if the head value is more than the tail value. If the queue is filled then the GPU waits for the CPU to process atleast one request. 1 Then the GPU increments the head to the next element. 2 The CPU thread continuously reads the element at the tail to see if there is a request from the GPU, and when there is a request, the CPU reads the request, zeros out the current element, and increments the tail. 3 Now the CPU thread will process the request. Below we explain the requests generated by primitives and how the CPU handle these requests.
@@ 4.2.1. PortChannel
Synchronization. 5 signal pushes a signal request in the queue. 6 7 The CPU processes this request by atomically incrementing the semaphore on the receiving GPU using functions, like ibv_atomic_add for IB. The wait primitive on the receiving GPU do not create a request for its CPU, therefore, the receiving CPU is idle, rather wait continuously looks for the semaphore to reach a expected value in a while loop. 8 wait returns after the semaphore is incremented.
Flush. Primitive flush pushes a flush request in the queue and the first thread of the GPU waits until the queue head is equal to or more than the queue tail. The CPU processes the flush request by waiting until all the previous data transfer and synchronization requests have been completed. For example, for IB we use ibv_poll_cq to get the status of all requests. After the flush request is complete, the GPU is free to re-write to the source buffer.
4.2.2. MemoryChannel
A MemoryChannel wraps data transfer methods that use thread-copy mechanism, i.e., directly use GPU threads for writing to peer GPUs memory.
Protocols. The channel provides two protocols, which define data transfer and synchronization technique to tune between low-latency and high-bandwidth: HB protocol provides a high-bandwidth but high-latency protocol, thus, is suitable for larger sizes, and LL protocol provides low-latency but low-bandwidth, thus, is suitable for smaller sizes. Both protocols achieves these properties by synchronizing on the suitable granularity of data.
@@ 4.2.2. MemoryChannel
(ii) In LL protocol, the receiving GPU synchronizes on fewer chunks of data being transferred and process the chunk as soon as it is received, thus, achieving lower latency. However, since the number of synchronization is proportional to the number of elements, this protocol provides lower-bandwidth. The LL protocol works as follows. The put primitive requires an extra integer flag value. For every N1N-1 elements written to the receiving GPU, put also writes the flag. The receiving GPU uses the read primitive, waits until the flag value at $N$ index of receiving buffer is set, and then reads and returns the N1N-1 elements. However, we cannot arbitrarily use any $N$ because GPUs follow a weak memory consistency model where writes to different memory locations by multiple threads can be performed in any order. Therefore, we restrict $N$ to number of elements written by a single instruction, i.e., 4, 8, and 16 bytes memory accesses instructions. The flag value is decided based on the algorithm, such that, all flag values are distinct.
Implementation. A MemoryChannel is initialized with a protocol. A protocol defines the implementation of data transfer and synchronization primitives. We now describe the implementation of primitives based on the protocol as follows:
Data Transfer. Primitive put reads data elements from source buffer and writes to destination buffer starting from different offsets. Similarly, the read primitive reads and returns data from the destination buffer, and the write primitive writes data including the flag for the LL protocol to the destination buffer. To maximize bandwidth for the HB protocol, both primitives use 16-bytes loads and stores. For the LL protocol, both primitives also takes a flag and uses 8-byte load and stores by default or user supplied vector length. Since both put primitives are called by multiple threads (all or first few threads of the kernel), we achieve maximum bandwidth.
@@
4.2.3. SwitchChannel
A SwitchChannel provides primitives for performing collective operations among GPUs. These operations usually require specialized hardware support. For example, NVIDIA NVLink 4.0 connects all H100 GPUs to a single NVSwitch and this NVSwitch can perform collective operations including reduce and broadcast. A SwitchChannel in MSCCL++ provides two primitives: reduce to add the corresponding elements of buffers residing on different GPUs and broadcast to send elements from a buffer on a GPU to all other GPUs. We now discuss the implementation of these primitives for NVSwitch, which provides in-network, switch-based aggregation and multicast capabilities using NVLink SHARP (NVLS) technology (NVIDIA, 2025c). However, we believe that these implementations can be generalized to other hardware in the future.

### v4

@@ 3. MSCCL++Overview
MSCCL++ Primitive API. This replaces NCCL’s primitives with a set of core communication functions that are called from the GPU kernels. This is paired with the bootstrapping API called from the host (CPU) side, which is used for connection setup between GPUs. The primitive interface is the base layer that is used to implement the other two higher-level interfaces. Programmers can also use the primitive interface directly to implement application-specific or hardware-specific optimizations in their own GPU kernels, including those kernels that also perform computation.
@@
MSCCL++ DSL API. Based on the primitive interface, we introduce a new DSL to write custom collective communication algorithms in a high-level language. This interface is aimed at users who want to specify communication algorithms or schedules and generate kernels that are optimized for their own workloads (application and hardware). The generated kernel uses the Primitive API by running the DSL Executor, so it may have small runtime overheads.
@@
MSCCL++ Collective API. We re-implement the NCCL API, including its bootstrapping API, as-is over MSCCL++. It is aimed at users with the least expertise; they can simply replace NCCL/RCCL with the MSCCL++ Collective library without changing their application code. Similar to NCCL/RCCL, this library may not provide the best algorithm for certain workloads. For AI inference workloads, we implement the best algorithms in our collective kernels using the MSCCL++ DSL.
4. MSCCL++Primitive Abstractions
1class PortChannel
2 void *src, *dst; // set during init
3 uint *semaphore; uint expectedVal;
4
5 void put(ulong dstOff, ulong srcOff, ulong sz);
6 void signal();
7 void wait();
8 void flush();
9
10template<Protocol protocol> // LL or HB protocol
11class MemoryChannel
12 void *src, *dst; // set during init
13 uint *semaphore; uint expectedVal;
14
15 void put(ulong dstOff, ulong srcOff, ulong sz,
16 uint tid, uint tids, uint flag=0);
17 template<typename T>
18 T read(ulong off, uint flag=0);
19 template<typename T>
20 void write(ulong off, T elem, uint flag=0);
21 void signal();
22 void wait();
23 void flush();
24
25class SwitchChannel
26 void *local, *multimem; // set during init
27
28 void reduce(ulong dstOff,ulong srcOff,ulong sz);
29 void broadcast(ulong dstOff,ulong srcOff,
30 ulong sz);
Figure 2. Channel abstractions in MSCCL++ Primitive API.
MSCCL++Primitive API provides the minimal hardware abstractions for GPU communication. These abstractions are close to the hardware to avoid performance overheads. At the same time, they hide the complexities of GPU/CPU/NIC orchestration, synchronization, and memory consistency issues to simplify the programmability.
@@ 18 T read(ulong off, uint flag=0);
The modes of data transfer differ across different interconnects. At its core, MSCCL++ provides the notion of a Channel to capture these modes of communication with basic synchronization primitives. Specifically, MSCCL++ provides (1) PortChannel for port-mapped I/O allowing GPUs to initiate data transfer through dedicated hardware I/O ports such as DMA engines or RDMA NICs, (2) MemoryChannel for memory-mapped I/O with direct peer-to-peer memory access across GPUs, and (3) SwitchChannel for switch-mapped I/O using multicasting memory access enabled by interconnect switches. Each of these channels exposes additional primitives specific to that channel.
An interconnect might support multiple modes of data transfer and hence can support multiple of these interfaces. For instance, NVLink supports all the three, while xGMI and PCIe only support port-mapped and memory-mapped I/O. Our intent is that these abstractions are general allowing the library implementer to quickly adapt to changing hardware capabilities. For instance, the current implementation of PortChannel for NVLink requires coordination with the CPU to initiate a DMA engine (cudaMemcpy), but can directly initiate it from the GPU if future hardware supports it.
@@
Port-mapped, memory-mapped, and switch-mapped I/O are complementary I/O methods in general computer architectures. PortChannel, MemoryChannel, and SwitchChannel are our abstractions to provide these I/O methods from inside GPU kernels. We believe that they are flexible enough to accommodate future hardware advances. For example, if future hardware interconnects support initiating DMA-copy from within a GPU kernel (Hwang et al., 2023b), the same PortChannel API can target the new hardware.
@@
Figure 2 shows three types of channels and its interface in MSCCL++. MSCCL++ also supports a few fused primitives (omitted in Figure 2 for brevity) that can reduce the overhead of API calls. For example, since a put is usually followed by a signal, we provide the put_with_signal function that performs both at once. The buffers and semaphores are allocated on the GPU memory before initialization of the channels. Next, we describe the three kinds of MSCCL++ channels.
4.1. PortChannel
A PortChannel implements primitives when data transfer is done over ports connected to GPU memory, such as cudaMemcpy for intra-node DMA or ibv_post_send for RDMA. Borrowing the term from MPI (Snir et al., 1998), writing data to the peer’s side is called put. To provide our communication abstraction as close as possible to hardware capabilities, we design put to be zero-copy (i.e., no intermediate buffers), one-sided (i.e., initiated by a peer without participation of the other peer), and asynchronous.
Figure 3 shows the semantics of the synchronization primitives: signal, wait, and flush. The put primitive transfers data from src0 buffer of GPU-0 to dst1 buffer of GPU-1. The signal is an asynchronous operation that is strictly ordered with respect to the previous transfer operations, such as put here. The matching wait primitive called by GPU-1 blocks till it receives GPU-0’s signal at which time GPU-1 is allowed to access the transferred data in dst1. GPU-0 calls the flush primitive to ensure that the transfer is complete from its perspective allowing it to reuse the src0 buffer.
Implementation. Since data transfer over a port currently requires the CPU to initiate the transfer, each channel creates its own CPU thread. The CPU thread reads data transfer or synchronization requests from a first-in-first-out request queue. The storage and tail of a request queue are allocated using cudaMallocManaged as both CPU and GPU can access it, while the head is on GPU and only accessed by GPU.
We now discuss the workflow of PortChannels using Figure 4. 0 When the GPU calls a primitive, such as put, the first participating thread of the GPU pushes the request to the queue by writing at the head. Before writing to the queue, the GPU checks if the queue is filled, i.e., if the head value is more than the tail value. If the queue is filled then the GPU waits for the CPU to process at least one request. 1 Then the GPU increments the head to the next element. 2 The CPU thread continuously reads the element at the tail to see if there is a request from the GPU, and when there is a request, the CPU reads the request, zeros out the current element, and increments the tail. 3 Now the CPU thread will process the request. Below we explain three types of requests and how the CPU handles them over InfiniBand (IB) as an example.
//async
put(src0, dst1, size)
//unsafe to reuse src0
signal() //async
flush() //sync
//safe to reuse src0
@@ 4.1. PortChannel
Synchronization. 5 signal pushes a signal request in the queue. 6 7 The CPU processes this request by atomically incrementing the semaphore on the receiving GPU using functions, like ibv_atomic_add for IB. The wait primitive on the receiving GPU does not create a request for its CPU, therefore, the receiving CPU is idle, rather wait continuously looks for the semaphore to reach an expected value in a while loop. 8 wait returns after the semaphore is incremented.
Flush. Primitive flush pushes a flush request in the queue and the first thread of the GPU waits until the queue tail reaches the position where the flush request is pushed to. The CPU processes the flush request by waiting until all the previous data transfer and synchronization requests have been completed. For example, for IB we use ibv_poll_cq to get the status of all requests. After the flush request is complete, the GPU is free to re-write to the source buffer.
4.2. MemoryChannel
A MemoryChannel wraps data transfer methods that use thread-copy mechanism, i.e., directly use GPU threads for writing to peer GPU’s memory.
Protocols. The channel provides two protocols, which define data transfer and synchronization technique to tune between low-latency and high-bandwidth: HB protocol provides a high-bandwidth but high-latency protocol, thus, is suitable for larger sizes, and LL protocol provides low-latency but low-bandwidth, thus, is suitable for smaller sizes. Both protocols achieve these properties by synchronizing on the suitable granularity of data.
@@ 4.2. MemoryChannel
(ii) In LL protocol, the receiving GPU synchronizes on fewer chunks of data being transferred and processes the chunk as soon as it is received, thus achieving lower latency. However, since the number of synchronization is proportional to the number of elements, this protocol provides lower-bandwidth. The LL protocol works as follows. The put primitive requires an extra integer flag value. For every N1N-1 elements written to the receiving GPU, put also writes the flag. The receiving GPU uses the read primitive, waits until the flag value at $N$ index of receiving buffer is set, and then reads and returns the N1N-1 elements. However, we cannot arbitrarily use any $N$ because GPUs follow a weak memory consistency model in which writes to different memory locations by multiple threads can be performed in any order. Therefore, we restrict $N$ to number of elements written by a single instruction, i.e., 4, 8, and 16 bytes memory accesses instructions. The flag value is decided based on the algorithm, such that, all flag values are distinct.
Data Transfer. put is a collective-thread operations that reads data elements from source buffer and writes to destination buffer starting from different offsets using a set of threads. Similarly, the read primitive reads and returns data from the destination buffer, and the write primitive writes data including the flag for the LL protocol to the destination buffer. To maximize bandwidth for the HB protocol, both primitives use 16-bytes loads and stores. For the LL protocol, both primitives also take a flag and use 8-byte loads and stores by default or user supplied vector length. Since both put primitives are called by multiple threads (all or first few threads of the kernel), we achieve maximum bandwidth.
@@
4.3. SwitchChannel
A SwitchChannel provides two primitives for performing collective operations among GPUs: reduce to add the corresponding elements of buffers residing on different GPUs and broadcast to send elements from a buffer on a GPU to all other GPUs. These operations usually require specialized hardware support. For example, NVIDIA NVLink 4.0 connects all H100 GPUs to a single NVSwitch and this NVSwitch can perform in-network, switch-based aggregation and multicast using NVLink SHARP (NVLS) technology (NVIDIA, 2025a). We now discuss the implementation of these primitives for NVSwitch, but we believe that these implementations can be generalized to other hardware in the future.